OpenStudy (anonymous):

solve the equation x/3-5=y for x A. x=3y-5 B.x=3y-15 C.x=3y+15 D.x=3y+5 could you help me on this one @hartnn sorry to bother you again

OpenStudy (ineedhelplz):

Step 1: Add 5 to both sides. \[\frac{ 1 }{ 3 } x−5+5=y+5\] \[\frac{ 1 }{ 3 }x=y + 5\]

OpenStudy (ineedhelplz):

Step 2: Divide both sides by 1/3. \[x=3y+15\] is the answer!

OpenStudy (anonymous):

youre amazingggg thank you so much for not just iving the answer but explaining how to get it

OpenStudy (ineedhelplz):

You're welcome :)